If one card is chosen from a standard pack, what is the probability that it is a diamond (♦) or a club (♣)?
Correct answer B. 26/52 = 1/2
Explanation
There are 13 diamonds and 13 clubs, and the suits are mutually exclusive, so P = 13/52 + 13/52 = 26/52 = 1/2.
Answer checked How we write and check questions
Discussion